OpenStudy (motherhen544):

Simplify. x^−5/x Write your answer with a positive exponent only.

OpenStudy (motherhen544):

\[\frac{ x ^{-5} }{ x }\]

OpenStudy (mathmale):

Unfortunately, this expression is ambiguous (subject to misinterpretation or interpretation in different ways. Can you think of any way in which to remove the ambiguity or ambiguities, using parentheses?

OpenStudy (motherhen544):

no this really confused me so i have no idea what to do

OpenStudy (mathmale):

Your \[\frac{ x ^{-5} }{ x }\] is so much better. Now, use rules of exponentiation to simplify, aiming for a positive exponent in the denominator.

OpenStudy (mathmale):

Actually, by presenting \[\frac{ x ^{-5} }{ x }\] you have remvoed all ambiguities. Perfect. Now, simplify your result.

OpenStudy (motherhen544):

but what does it mean by simplify it

OpenStudy (motherhen544):

1\[\frac{ 1 }{ x^{6}}\]

OpenStudy (motherhen544):

is that right?

OpenStudy (mathmale):

Before simplification: You have a neg exponent and a positive one. After simplification: You have combined those two exponents and come up with a positive one that goes into the denominator. Yes, your final answer is correct.